OpenStudy (steve816):

Please I need help. A 1-kg block of wood is dropped from a height 2.26 meters above a resting 0.05-kg ball. Shortly thereafter, the ball is shot horizontally at a velocity of 2.91 m/s. The wood is 9.94 meters away from the ball in the horizontal direction. Determine the distance (in meters) the wood falls from its original position until impact? Ignore air resistance.

OpenStudy (518nad):

solved this for u earlier

OpenStudy (steve816):

a=9.81m/s v=9.81 * t d=9.81/2*t^2 d=9.81/2*(9.94/2.91)^2

OpenStudy (steve816):

This is what you said earlier, and I'm still confused. I put your answer in my calculator and got 57.23, how is that a possible answer if the wood is dropped from 2.26 meters above?

OpenStudy (sooobored):



OpenStudy (sooobored):

such a weird problem

OpenStudy (steve816):

I know right? Blame my physics teacher and webassign. Do you think you can help @sooobored ?

OpenStudy (sooobored):

the shortly thereafter is... not very nice and vague

OpenStudy (sooobored):

ill give it a shot

OpenStudy (sooobored):

block of wood- at no initial velocity, and has initial height which i will define as h+2.26 ball- initial height of h, initial x velocity of 2.91m/s, 9.94m away from block in horizontal direction

OpenStudy (sooobored):